OpenStudy (anonymous):

how to decrease $60 by 20%

OpenStudy (gudden):

@zari3200 ... multiply 60 with 20 / 100 and then subtract this value from 60 and you would get the answer..

OpenStudy (kc_kennylau):

A number n decreased by x% = n(1-x%) A number n increased by x% = n(1+x%)

OpenStudy (kc_kennylau):

Percentage change from A to B = (B-A)/A * 100%